A book store can purchase several calculators for a total cost of $120. If each calculator cost $1 less, the bookstore could purchase 10 additional calculators at the same total cost. How many calculators can be purchased at the regular price?
Answer by TimothyLamb(4379)

---
x = number of calculators
y = price
---
xy = 120
(x + 10)(y - 1) = 120
---
y = 120/x
xy - x + 10y - 10 = 120
x120/x - x + 10*120/x - 10 = 120
-x + 1200/x - 10 = 0
-xx + 1200 - 10x = 0
-xx - 10x + 1200 = 0
---
the above quadratic equation is in standard form, with a=-1, b=-10, and c=1200
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
-1 -10 1200
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
this quadratic has two real roots (the x-intercepts), which are:
x = -40
x = 30
---
a negative number of calculators doesn't make sense for this problem, so use the positive root:
---
answer:
x = number of calculators = 30
y = price = $4
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php